### v1

# FaithSCAN: Model-Driven Single-Pass Hallucination Detection for Faithful Visual Question Answering

Abstract
Faithfulness hallucinations in VQA occur when vision-language models produce fluent yet visually ungrounded answers, severely undermining their reliability in safety-critical applications. Existing detection methods mainly fall into two categories: external verification approaches relying on auxiliary models or knowledge bases, and uncertainty-driven approaches using repeated sampling or uncertainty estimates. The former suffer from high computational overhead and are limited by external resource quality, while the latter capture only limited facets of model uncertainty and fail to sufficiently explore the rich internal signals associated with the diverse failure modes. Both paradigms thus have inherent limitations in efficiency, robustness, and detection performance. To address these challenges, we propose FaithSCAN: a lightweight network that detects hallucinations by exploiting rich internal signals of VLMs, including token-level decoding uncertainty, intermediate visual representations, and cross-modal alignment features. These signals are fused via branch-wise evidence encoding and uncertainty-aware attention. We also extend the LLM-as-a-Judge paradigm to VQA hallucination and propose a low-cost strategy to automatically generate model-dependent supervision signals, enabling supervised training without costly human labels while maintaining high detection accuracy. Experiments on multiple VQA benchmarks show that FaithSCAN significantly outperforms existing methods in both effectiveness and efficiency. In-depth analysis shows hallucinations arise from systematic internal state variations in visual perception, cross-modal reasoning, and language decoding. Different internal signals provide complementary diagnostic cues, and hallucination patterns vary across VLM architectures, offering new insights into the underlying causes of multimodal hallucinations.
I Introduction
Visual Question Answering (VQA) is a representative image-to-text (I2T) task in which a model generates natural language answers given an image and a corresponding question [huang2025survey]. Recent advances in multimodal large language models (MLLMs) and vision-language models (VLMs) have enabled impressive integration of visual perception and language understanding, supporting increasingly complex open-ended VQA tasks [ging2024openended, Yu2024TowardsOV]. Despite impressive performance, they frequently generate fluent yet ungrounded or incorrect responses, a phenomenon known as hallucination [ji2023survey, farquhar2024detecting, bai2024hallucination], which in VQA settings manifests as responses misaligned with the input visual evidence or contradictory to factual knowledge [liu2024survey, hu2024mitigating]. Among them, faithfulness hallucinations [huang2025survey], where answers violate visual grounding, are particularly critical in VQA, since reliable factual verification depends on the model first attending to and correctly interpreting the image [li-etal-2023-evaluating]. Consequently, detecting faithfulness hallucinations is a necessary step for ensuring answer reliability and user trust, motivating the development of effective detection methods that can identify unfaithful answers (exmaples illustrated in Fig. 1).
Motivated by the importance of faithfulness hallucination detection in VQA, a variety of detection methods have been proposed in recent years. Existing approaches can be broadly grouped into two paradigms: external verification-based methods and uncertainty-driven methods [bai2024hallucination, chen2025survey]. External verification approaches detect hallucinations by explicitly checking cross-modal consistency between the generated answer and the visual input, often relying on auxiliary components such as object detectors, VQA models, scene graphs, or large language model (LLM) judges [li-etal-2023-evaluating, jing2024faithscore, gu2024survey]. By decomposing answers into atomic claims and validating them against visual evidence, these methods are capable of identifying fine-grained inconsistencies. However, their reliance on additional pretrained models and multi-stage pipelines leads to high computational overhead and makes the final performance sensitive to error propagation from external components. Moreover, such methods operate largely independently of the internal reasoning process of the target VLM, limiting their scalability and integration in real-world deployment [jing2024faithscore].
Figure 1: Illustration of three representative VQA scenarios: (a) object existence questions, (b) relationships between two objects within a single image, and (c) reasoning with a false-premise question. Modern VLM-based QA systems perform reliably on simple questions, but often produce unfaithful or hallucinated answers in complex scenarios.
In contrast, uncertainty-driven methods approach hallucination detection from the perspective of model confidence. Empirical studies in both language models and vision–language models have shown that hallucinated outputs are often associated with elevated predictive uncertainty [huang2025survey, liu2024survey]. Representative approaches, such as semantic entropy (SE) [farquhar2024detecting] and likelihood-based measures [farquhar2024detecting, zhang2024vl, ding2024hallu, tong2025semantic], estimate hallucination likelihood through sampling model outputs under different decoding strategies, prompt reformulations, or visual perturbations, followed by entropy or distributional analysis. Recent work enhances uncertainty estimation by incorporating visual cues, such as vision-amplified semantic entropy [liao2025vision] and dense geometric entropy [gautam2025hedge]. In addition, methods leveraging internal signals like attention patterns and intermediate embeddings are emerging. Approaches such as DHCP [zhang2025dhcp] and OPERA [huang2024opera] focus on attention dynamics to detect over-reliance on unverified features, while HallE-Control [zhai2023halle] modifies embeddings to suppress unreliable visual details. These methods highlight the value of integrating internal uncertainty signals.
Despite their effectiveness, existing uncertainty-based methods suffer from two key limitations. First, they rely on repeated sampling and generation-level statistics, resulting in high inference cost and unstable behavior. This dependence on stochastic decoding and prompt configurations significantly limits their practicality in latency-sensitive or large-scale VQA settings. Second, they fail to jointly leverage the heterogeneous uncertainty signals inherently encoded in modern VLMs. Uncertainty arising from language decoding, visual perception, and vision–language interaction is often treated independently or partially ignored, rather than being modeled in a unified manner. As a result, the complementary information carried by different internal uncertainty sources remains largely unused for hallucination detection.
To tackle these limitations, we propose FaithSCAN, a lightweight faithfulness hallucination detection network for VQA. FaithSCAN treats hallucination detection as a supervised learning problem, where multiple sources of internal uncertainty are extracted from a single forward pass of a VLM. Specifically, it leverages three complementary uncertainty signals: (i) token-level predictive uncertainty from language decoding, (ii) intermediate visual representations reflecting the model’s perception of the image, and (iii) cross-modal aligned representations that project visual information into the textual semantic space. These heterogeneous signals are processed via Branch-wise Evidence Encoding and aggregated using Uncertainty-Aware Attention, enabling the detector to capture hallucination patterns across multiple reasoning pathways. Supervision is provided through a model-driven Visual-NLI procedure, extending the LLM-as-a-Judge paradigm to VQA, and is further validated by small-scale human verification, ensuring label reliability without requiring large-scale manual annotation.
Our contributions are summarized as follows:
  • We introduce FaithSCAN, a novel hallucination detection architecture that leverages internal uncertainty signals from a single VLM forward pass.
  • FaithSCAN jointly models uncertainty from language decoding, visual perception, and cross-modal interaction through branch-wise encoding and uncertainty-aware aggregation, linking internal signals to hallucinations.
  • We propose a model-aware supervision strategy by extending the LLM-as-a-Judge paradigm to multimodal hallucination detection in VQA, where hallucination labels are directly aligned with the model’s internal reasoning processes through a Visual-NLI framework.
  • We conduct comprehensive experiments on four major VQA datasets, addressing various hallucination types, and demonstrate FaithSCAN’s effectiveness across multiple research questions.
The remainder of this paper is organized as follows. Section II reviews related work. Section III formally defines the problem setting and introduces the key concepts used throughout this study. Section IV presents the proposed methodology. Section V describes the experimental setup, including datasets, evaluation metrics, and implementation details. Section VI reports experimental results and analysis. Finally, Section VII concludes the paper.
II Related Work
II-A Hallucination Detection in Visual Question Answering
Faithfulness hallucination detection is critical in VQA, where generated answers can be fluent and confident yet inconsistent with visual evidence [chen2025survey]. A variety of detection methods have been proposed in recent years, which can be broadly categorized into external verification-based and internal signal- or uncertainty-driven approaches. External verification methods explicitly check cross-modal consistency between answers and visual inputs, often using object detectors, VQA models, scene graphs, or LLM evaluators [li-etal-2023-evaluating, jing2024faithscore, gu2024survey]. By decomposing answers into atomic claims and validating them against visual evidence, they enable fine-grained hallucination detection. However, these methods rely on multi-stage pipelines and pretrained components, incurring high computational cost and sensitivity to error propagation, while largely ignoring the VLM’s internal reasoning process [jing2024faithscore]. Internal signal- or uncertainty-driven methods, in contrast, exploit the model’s outputs, attention patterns, and intermediate embeddings to detect hallucinations, offering a more direct and efficient view into the generation process [liu2025paying, jiang2025devils, ghosh2025visual].
II-B Uncertainty-Based Hallucination Detection
Model-internal uncertainty has emerged as an effective and relatively inexpensive cue for hallucination detection. Both language models and VLMs exhibit increased predictive uncertainty when hallucinations occur [zhang2023enhancing, quevedo2024detecting, aryal2025howard, gu2024survey]. Methods such as semantic entropy (SE) [farquhar2024detecting] and likelihood-based measures [ma2025semantic, zhao2025sese, jiang2025interpreting] quantify uncertainty by aggregating variations across multiple decoding outputs. In VQA, extensions incorporate visual perturbations or multimodal semantic clustering, e.g., vision-amplified semantic entropy [liao2025vision] and dense geometric entropy [gautam2025hedge], which estimate hallucination likelihood by comparing semantic distributions under clean and perturbed visual inputs. These methods demonstrate that uncertainty signals can effectively capture hallucinations, including in safety-critical domains such as medical VQA [liao2025vision]. Nevertheless, most approaches rely on repeated sampling or perturbations to obtain reliable estimates, which increases inference cost and can be sensitive to stochastic decoding behavior.
II-C Sources and Exploitation of Internal Uncertainty Signals
Hallucinations in VQA are not only reflected in the magnitude of predictive uncertainty but also in structured patterns across attention, intermediate representations, and logits. Early studies leveraged attention dynamics to detect hallucinations, e.g., DHCP [zhang2025dhcp] captures cross-modal attention patterns, showing consistent differences between hallucinated and faithful responses. OPERA [huang2024opera] demonstrates that hallucinations correlate with over-reliance on sparse summary tokens in self-attention, which can be mitigated through dynamic reweighting. Feature-based methods, such as HallE-Control [zhai2023halle], reduce hallucinations by projecting intermediate embeddings to suppress reliance on unverifiable visual details. Embedding- or logit-based approaches, e.g., PROJECTAWAY [jiang2025interpreting] and ContextualLens [phukan2025beyond], exploit token-level predictions and context-aware embeddings to localize and mitigate hallucinations at fine-grained object or patch levels. Despite their effectiveness, these methods are often limited by partial access to internal signals or simplified modeling of cross-modal interactions. Together, they indicate that hallucination-related cues are encoded across heterogeneous internal signals, highlighting the feasibility and necessity of fully exploring and exploiting them for reliable detection.
More recently, HaluNet [tong2025halunet] studies efficient hallucination detection in text-based question answering using internal uncertainty signals. Our work instead focuses on faithfulness hallucinations in VQA and extends this paradigm through explicit fusion of multimodal internal evidence. In addition, we provide a more systematic design and empirical validation of this approach in the multimodal setting.
III Preliminaries
III-A Visual Question Answering
VQA aims to generate a natural language answer conditioned on an input image and a textual question. Formally, let:
$I\in\mathcal{I},\quad q\in\mathcal{Q},$ (1)
denote the input image and question, respectively, where $\mathcal{I}$ and $\mathcal{Q}$ denote the spaces of images and textual questions. Given $(I,q)$, a VLM parameterized by $\theta$ produces an answer:
$\hat{a}=f_{\theta}(I,q),\quad\hat{a}\in\mathcal{A},$ (2)
where $\mathcal{A}$ denotes the space of textual responses.
In this work, we focus on modern VLMs that generate answers in an autoregressive manner. Specifically, all three VLM architectures studied in this paper (see Section III-C) follow this paradigm, modeling the conditional distribution as:
$p_{\theta}(a\mid I,q)=\prod_{t=1}^{T}p_{\theta}(a_{t}\mid a_{<t},I,q),$ (3)
where $a$ denotes the output token sequence.
III-B Faithfulness Hallucination Detection in VQA
We focus on faithfulness hallucinations [huang2025survey], where the generated answer is inconsistent with the visual evidence present in the input image. Given a VQA instance:
[…]
the goal of hallucination detection is to determine whether the answer $\hat{a}$ is visually grounded in $I$.
We formulate faithfulness hallucination detection as a supervised binary classification problem. Specifically, we aim to learn a classifier:
$g_{\phi}:(I,q,\hat{a})\rightarrow y,\quad y\in\{0,1\},$ (5)
where $y=1$ indicates a faithfulness hallucination and $y=0$ denotes a faithful answer.
Unlike generation-based uncertainty estimation methods, our learning objective is not to approximate the full output distribution $p_{\theta}(a\mid I,q)$, but to directly model:
[…]
using uncertainty signals and representations internally encoded in the VLM during a single forward pass. This formulation enables efficient hallucination detection without repeated sampling, external verification models, or additional knowledge sources.
TABLE I: Explicit mapping from observed hallucination phenomena to failure pathways in the multimodal reasoning process.
Observed Hallucination Phenomenon Typical Manifestation in VQA Primary Failure Pathway
Object Hallucination Mentioning non-existent objects or attributes not supported by the image Perceptual Failure: inaccurate or incomplete visual entity representation
Relational Hallucination Incorrect spatial, logical, or attribute relations between visual entities Perceptual $\rightarrow$ Reasoning Failure: unreliable visual grounding propagates to relational inference
Scene Hallucination Misinterpreting global scene context or high-level visual semantics Perceptual $\rightarrow$ Reasoning Failure: failure in compositional reasoning and cross-modal semantic alignment
False-Premise Hallucination Answering questions with invalid or contradictory premises without rejection Interaction-Level Failure: incorrect grounding of the question–answer interaction
III-C Vision–Language Model Architectures
Existing VLMs differ in architectural design, but most share three functional components: a vision encoder, a language model, and a cross-modal alignment mechanism. We consider three representative models that cover distinct alignment paradigms.
III-C1 InstructBLIP
InstructBLIP [dai2023instructblip] employs a query-based cross-modal alignment strategy. A frozen vision encoder first extracts dense visual features from the image. These features are then queried by a lightweight Query Transformer (Q-Former), producing a compact set of visual query embeddings. The embeddings are projected into the input space of a Flan-T5 language model, enabling instruction-following VQA without fully fine-tuning the vision backbone or language model.
This design emphasizes compact visual abstraction and controlled vision–language interaction, but limits fine-grained token-level coupling between visual and textual modalities.
III-C2 LLaVA-NeXT
LLaVA-NeXT adopts the same projection-based multimodal alignment paradigm as the original LLaVA [liu2023visual]. Its vision encoder (vision tower) extracts high-dimensional visual features, which are projected by a learnable multimodal layer and concatenated with text token embeddings. The combined embeddings are fed into a decoder-only LLaMA model, enabling cross-modal reasoning while treating visual features as additional tokens aligned to the language space. We use LLaVA-NeXT in this work as a stronger instantiation of the LLaVA architecture.
III-C3 Qwen-VL
Qwen3-VL[Qwen3-VL] adopts a unified multimodal transformer architecture. Visual features are embedded and merged into a sequence of multimodal tokens, which are processed alongside textual tokens across multiple transformer layers. Unlike projection-based models, visual and textual representations interact throughout the network, allowing richer semantic alignment. This tight integration improves cross-modal reasoning but also complicates uncertainty propagation, which is critical for hallucination detection.
III-D Implications for Hallucination Detection
Despite architectural differences, all three VLMs encode rich uncertainty information in their internal representations, including token-level predictive uncertainty, visual representation reliability, and cross-modal semantic alignment. These shared properties motivate a general hallucination detection approach that leverages internal representations and uncertainty signals, rather than relying on generation-level sampling or external verification.
IV Methodology
IV-A Method Overview
In this work, we propose a single-pass, model-driven framework for multimodal hallucination detection in VQA (Fig. 2). Unlike prior uncertainty-based methods that rely on repeated sampling or external verification, our approach operates entirely within a single forward pass of a VLM, leveraging its internal representations as supervisory signals. Hallucination detection is formulated as a supervised classification problem informed by multiple complementary sources of uncertainty: (i) token-level predictive uncertainty during answer generation, (ii) intermediate visual representations encoding the model’s perception, and (iii) visually grounded representations aligned with the textual semantic space. Being fully model-driven, the method does not require external estimators or auxiliary generative processes, making it compatible with diverse VLM architectures such as query-based (InstructBLIP), projection-based (LLaVA), and tightly coupled multimodal (Qwen3-VL) models. By explicitly learning the mapping between observed hallucination phenomena, their manifestations in VQA, failure pathways of multimodal reasoning, and internal model states (Table I), our approach provides a principled and efficient solution for hallucination detection without expensive inference cost.
Figure 2: Overview of FaithSCAN for single-pass hallucination detection in VQA. (a) Workflow of our method: multiple types of internal uncertainty signals are extracted from a frozen VLM in a single forward pass and combined with model-driven supervision to train FaithSCAN. (b) Core components of FaithSCAN, including the construction of supervision signals and the model architecture.
IV-B Why Supervised Hallucination Detection
Most uncertainty-based hallucination detection methods [farquhar2024detecting, zhang2024vl, ma2025semantic, tong2025semantic, zhao2025sese] aim to estimate or calibrate uncertainty, yet converting these scores into decisions still requires carefully chosen thresholds, limiting practical utility. By framing detection as a supervised decision-making problem, we directly map internal uncertainty signals to hallucination labels, producing actionable outputs and simplifying deployment. As described in Section IV-E, this approach learns from internal signals rather than calibrating model uncertainty, reducing learning difficulty: learning focuses on extracting useful internal signals rather than output-space confidence modeling.
IV-C Low-Cost Supervision via Model-Driven Labeling
A central challenge in supervised hallucination detection is the acquisition of reliable supervision at scale. While manual annotation of multimodal hallucinations is prohibitively expensive, many existing benchmarks provide only answer-level correctness or predefined hallucination references [goyal2017vqa, li-etal-2023-evaluating, wang2024haloquest, park2025halloc]. Similar research has tackled this challenge [chen2024mllm], however, hallucination annotations generated by these methods are typically model-agnostic, focusing on generic failure cases without considering that hallucinations are inherently model-dependent phenomena. Since hallucinations arise from specific generation behaviors, inductive biases, and cross-modal alignment failures of the underlying vision–language model, decoupling hallucination labels from the base model limits their practical utility for detection.
In contrast, we argue that reliable supervision for hallucination detection should be model-aware, grounded in the actual failure modes exhibited during inference. To this end, we propose a low-cost, model-driven supervision strategy inspired by the LLM-as-a-Judge paradigm [gu2024survey, chen2024mllm], and extend it to the multimodal hallucination setting, as shown in Fig. 2b. The core idea is to leverage a strong vision–language model as a semantic judge to evaluate whether a candidate answer is faithfully grounded in the visual input and consistent with the question–answer pair, with the image serving as explicit evidence.
Unlike previous work, our approach incorporates a novel model-driven framework by explicitly integrating visual grounding constraints through a structured Visual-NLI (Visual Natural Language Inference) protocol. This allows hallucination labels to be tightly coupled with the model’s internal state and generation behavior. Importantly, our goal is not to introduce an additional black-box oracle for hallucination detection, but rather to design the supervision process to remain fundamentally tied to the underlying model’s inference mechanics. Specifically, the supervision signal is derived from two complementary aspects: (i) hallucination labels are conditioned on semantic consistency judgments that align with the same image–question context used by the target model, directly linking the labels to the model’s internal reasoning processes; and (ii) the supervision process is coupled with the model’s generation behavior, ensuring that the generated hallucination labels reflect the specific failure modes exhibited by the model during inference.
To make this supervision signal reliable, we employ two targeted strategies. The following sections describe these components in detail.
IV-C1 Answer-Level Model-Driven Hallucination Labels with Gold References
We consider VQA datasets that provide a reference answer $a^{\ast}$ for each image–question pair $(I,q)$. Given a VLM $f_{\theta}$, we first obtain its most likely answer $\hat{a}$ under a low-temperature decoding regime. The objective is to determine whether $\hat{a}$ is faithful to the visual evidence and semantically consistent with the reference answer.
To this end, we formulate hallucination labeling as a visual natural language inference (VNLI) problem. Specifically, we construct the triplet $(I,q,a^{\ast})$ as the premise and treat the model-generated answer $\hat{a}$ as the hypothesis. A pretrained vision–language judge $\mathcal{J}$ is then queried to assess the semantic relationship between the premise and the hypothesis, conditioned on the image.
Formally, the judge outputs a categorical probability distribution:
$\mathbf{p}=\mathcal{J}(I,q,a^{\ast},\hat{a})\in\Delta^{3},$ (7)
where $\mathbf{p}=(p_{\text{ent}},p_{\text{con}},p_{\text{unc}})$ corresponds to the probabilities of entailment, contradiction, and uncertainty, respectively. The three judgment categories are defined as follows:
  • Entailment: the hypothesis $\hat{a}$ is semantically consistent with the reference answer $a^{\ast}$, and any additional details are clearly supported by the image.
  • Contradiction: the hypothesis conflicts with the reference answer or describes visual attributes that are inconsistent with the image.
  • Uncertainty: the image does not provide sufficient evidence to verify the hypothesis, or the hypothesis introduces details that are unsupported but not directly contradicted.
To improve robustness against stochasticity in the judge model, we perform $R$ independent judging rounds and average the predicted probabilities:
$\bar{\mathbf{p}}=\frac{1}{R}\sum_{r=1}^{R}\mathbf{p}^{(r)}.$ (8)
We then define the hallucination probability as the aggregated mass of non-entailment outcomes:
$p_{\text{hall}}=\bar{p}_{\text{con}}+\bar{p}_{\text{unc}}.$ (9)
The final answer-level hallucination label is determined by:
$y_{\text{hall}}=\mathbb{I}\!\left(p_{\text{hall}}>\bar{p}_{\text{ent}}\right),$ (10)
where $\mathbb{I}(\cdot)$ denotes the indicator function.
In this work, we treat both contradiction and uncertainty as indicators of hallucination, as they reflect failures of visual grounding or over-commitment beyond the available visual evidence.
We emphasize that the judge model is only used during dataset construction. At inference time, the proposed hallucination detector operates independently, relying solely on the internal representations of the target VLM.
IV-C2 Human Verification via Small-Scale Multi-Round Sampling
While the proposed model-driven labeling strategy enables scalable supervision, it is still essential to assess the reliability of the automatically constructed hallucination labels. To this end, we introduce a small-scale human verification protocol that serves as a quality control mechanism rather than a primary annotation source.
Concretely, we randomly sample a subset of VQA instances (100 samples per data set) and inspect the most-likely answer generated by the target VLM under low-temperature decoding. For each sampled instance, human annotators are presented with the image, question, gold answer, and model-generated answer, and must determine whether the model answer is a hallucination—defined as factually incorrect or unsupported by the visual evidence. To avoid confounding factors, samples with ambiguous questions or problematic ground-truth references are explicitly marked and excluded from subsequent agreement analysis.
Importantly, human annotation is not used to replace model-driven labels. Instead, it is employed to validate whether the hallucination signals produced by the Visual-NLI judge are consistent with human judgment when applied to the same model outputs. Let $y^{\text{human}}\in\{0,1\}$ denote the human label indicating non-hallucination or hallucination, and $y^{\text{model}}\in\{0,1\}$ denote the corresponding binary label derived from the model-driven NLI judgment. We evaluate the agreement between $y^{\text{human}}$ and $y^{\text{model}}$ using multiple complementary metrics, including raw agreement ratio, Cohen’s $\kappa$, and Matthews correlation coefficient (MCC); detailed analyses and discussions are provided in Section VI-B.
This verification process serves two purposes. First, it empirically assesses the reliability of the model-driven supervision signal without requiring large-scale manual annotation. Second, it provides evidence that hallucination labels constructed via Visual-NLI judgments are not arbitrary artifacts of the judge model, but are aligned with human perception of visual faithfulness. As such, the human verification step supports the use of model-driven labels as effective supervision for training hallucination detectors.
IV-D Uncertainty-Related Internal Signals in Vision–Language Models
In this subsection, we describe three categories of internal signals that are accessible in mainstream VLMs and serve as the foundation of our hallucination detection method. Our approach is based on the observation that faithfulness hallucinations in VLMs are closely tied to the model’s internal uncertainty during inference. Rather than introducing external supervision or repeated sampling, we directly extract uncertainty-related signals from intermediate representations produced in a single forward pass of the VLM $f_{\theta}$ (Fig. 2a). Importantly, these signals are model-driven and architecture-dependent: their concrete instantiation depends on the design of a given VLM, while their functional roles can be unified under a common formalization.
IV-D1 Problem Setting
Given an image–question pair $(I,q)$, the VLM generates an answer sequence $\hat{a}=(a_{1},\ldots,a_{T})$ autoregressively according to the conditional distribution $p_{\theta}(\hat{a}\mid I,q)=\prod_{t=1}^{T}p_{\theta}(a_{t}\mid a_{<t},I,q)$. During this generation process, the model produces a collection of intermediate representations spanning visual encoding, cross-modal alignment, and language decoding. We extract uncertainty-related signals from these internal representations without modifying the model parameters $\theta$ or the inference procedure.
IV-D2 Token-Level Generation Signals
At each decoding step $t$, the language model outputs a predictive distribution
$p_{\theta}(\cdot\mid a_{<t},I,q).$ (11)
From this distribution and the associated decoder hidden states, we extract three token-level signals: the log-likelihood of the generated token,
$\ell_{t}=\log p_{\theta}(a_{t}\mid a_{<t},I,q),$ (12)
the entropy of the predictive distribution,
$H_{t}=-\sum_{v\in\mathcal{V}}p_{\theta}(v\mid a_{<t},I,q)\log p_{\theta}(v\mid a_{<t},I,q),$ (13)
and the hidden embedding $\mathbf{h}_{t}\in\mathbb{R}^{d_{h}}$ of the generated token from a selected decoder layer. Here, $d_{h}$ denotes the dimensionality of the decoder hidden states. While $\ell_{t}$ and $H_{t}$ provide scalar measures of predictive confidence, the hidden embedding $\mathbf{h}_{t}$ is treated as a high-dimensional representation that implicitly encodes semantic uncertainty.
IV-D3 Visual Semantic Representations
Prior to cross-modal interaction, the vision encoder extracts visual features from the input image $I$. We denote the resulting set of patch-level visual embeddings as
$\mathbf{V}^{\mathrm{raw}}=\{\mathbf{v}_{i}\}_{i=1}^{N},\qquad\mathbf{v}_{i}\in\mathbb{R}^{d_{v}},$ (14)
where $N$ is the number of visual patches and $d_{v}$ denotes the dimensionality of the visual embedding space. Each $\mathbf{v}_{i}$ corresponds to a localized image region. In practice, $\mathbf{V}^{\mathrm{raw}}$ is obtained from the final hidden states of the vision backbone before any language conditioning or multimodal fusion. This representation reflects the model’s internal visual perception state and serves as a latent visual semantic space that may encode uncertainty in visual understanding.
IV-D4 Cross-Modal Aligned Representations
Modern VLMs employ an explicit alignment module to project visual representations into a language-compatible semantic space. We denote the resulting aligned visual tokens as:
$\mathbf{V}^{\mathrm{align}}=\{\tilde{\mathbf{v}}_{j}\}_{j=1}^{M},\qquad\tilde{\mathbf{v}}_{j}\in\mathbb{R}^{d_{a}},$ (15)
where $M$ is the number of aligned visual tokens and $d_{a}$ denotes the dimensionality of the aligned multimodal embedding space. These representations are produced by a model-specific alignment function:
$\mathbf{V}^{\mathrm{align}}=\mathcal{T}_{\theta}(\mathbf{V}^{\mathrm{raw}},q),$ (16)
which may incorporate cross-attention, projection layers, or query-dependent pooling mechanisms. The aligned visual tokens directly participate in the autoregressive decoding of $\hat{a}$ and therefore reflect visual information as mediated by vision–language interaction.
IV-D5 Unified View of Uncertainty Signals
Although the concrete extraction of the above signals depends on the architecture of the VLM, they can be jointly formalized as a set of internal uncertainty-related representations produced during the generation of $\hat{a}$, namely:
$\mathcal{S}_{\theta}(I,q,\hat{a})=\Big(\{(\ell_{t},H_{t},\mathbf{h}_{t})\}_{t=1}^{T},\mathbf{V}^{\mathrm{raw}},\mathbf{V}^{\mathrm{align}}\Big),$ (17)
which collectively capture uncertainty arising from language decoding, visual semantic perception, and vision–language interaction, and constitute a model-internal, single-pass characterization of uncertainty. In the next subsection, we describe how these heterogeneous signals are jointly modeled to produce a hallucination confidence score.
IV-D6 Unified View of Uncertainty Signals
Although the concrete extraction of the above signals is tightly coupled with the architecture of a given VLM, their roles are conceptually unified. Token-level generation signals capture uncertainty in language decoding, visual semantic representations reflect uncertainty in visual understanding, and aligned multimodal representations encode uncertainty arising from vision–language interaction.
IV-E Model Architecture
We propose FaithSCAN, a lightweight supervised hallucination detection network that operates on internal uncertainty signals extracted from a frozen VLM, as illustrated in Fig. 2b. FaithSCAN aggregates heterogeneous uncertainty evidence at the decision level and outputs a scalar hallucination score in a single forward pass. Notably, the detector is model-dependent: the dimensionality, number, and extraction locations of its inputs are determined by the architecture of the underlying VLM, enabling adaptation to model-specific uncertainty patterns without modifying the base model.
IV-E1 Input Formulation
Given the frozen VLM $f_{\theta}$, an image–question pair $(I,Q)$ and the model-generated answer $\hat{a}$, we extract $K$ internal uncertainty signals:
$\mathcal{E}=\{\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{K}\},\qquad\mathbf{X}_{k}\in\mathbb{R}^{L_{k}\times D_{k}}.$ (18)
Each $\mathbf{X}_{k}$ is a sequence-valued feature (e.g., token-level statistics, hidden states, or vision–language alignment features) obtained from different components of $f_{\theta}$. These sources may vary in modality, sequence length, and feature dimension, and are treated independently without assuming a shared feature space.
IV-E2 Branch-wise Evidence Encoding
Each uncertainty source $\mathbf{X}_{k}$ is processed by an independent branch encoder $f_{k}(\cdot)$ that compresses a variable-length sequence into a fixed-dimensional embedding:
$\mathbf{h}_{k}=f_{k}(\mathbf{X}_{k}),\qquad\mathbf{h}_{k}\in\mathbb{R}^{d},$ (19)
where $d=64$ in all experiments.
All branch encoders are implemented as lightweight sequence compression modules rather than deep semantic encoders. Specifically, we consider linear projections with global average pooling, trainable sequence compressors (Linear + LayerNorm + ReLU + mean pooling), and convolutional encoders with adaptive pooling [kim2014convolutional], where one-dimensional convolutions are applied with kernel size $3$ and stride $1$. Despite architectural differences, all variants share the same objective: to summarize the reliability and stability of internal signals while limiting representational capacity to reduce overfitting. We further analyze the generalization capability under distribution shifts in Section VI-A2, while the effectiveness of architectural variants is examined in Section VI-C2.
IV-E3 Uncertainty-aware Attention Fusion
Let $\mathbf{H}=[\mathbf{h}_{1},\ldots,\mathbf{h}_{K}]\in\mathbb{R}^{K\times d}$ denote the set of branch-wise evidence embeddings. To aggregate heterogeneous uncertainty sources, FaithSCAN employs a cross-branch attention mechanism:
$\alpha_{k}=\frac{\exp\!\left(\mathbf{w}_{a}^{\top}\tanh\!\left(\mathbf{W}_{a}\mathbf{h}_{k}\right)\right)}{\sum_{j=1}^{K}\exp\!\left(\mathbf{w}_{a}^{\top}\tanh\!\left(\mathbf{W}_{a}\mathbf{h}_{j}\right)\right)},\qquad\mathbf{h}=\sum_{k=1}^{K}\alpha_{k}\mathbf{h}_{k}.$ (20)
To further enhance adaptivity, we introduce a gated residual fusion:
$\mathbf{g}=\sigma\!\left(\mathbf{W}_{g}\mathbf{h}\right),\qquad\widetilde{\mathbf{h}}=\mathbf{h}\odot\mathbf{g}+\mathbf{h},$ (21)
where $\sigma(\cdot)$ denotes the sigmoid function. This formulation enables selective modulation of aggregated uncertainty evidence while preserving the original representation through the residual path.
IV-E4 Hallucination Prediction
The fused representation $\widetilde{\mathbf{h}}$ is mapped to a hallucination probability via:
$p=\sigma\!\left(\mathbf{w}^{\top}\widetilde{\mathbf{h}}+b\right),$ (22)
where $p\in(0,1)$ denotes the likelihood that the generated answer is hallucinated. FaithSCAN is trained end-to-end using binary cross-entropy loss with model-driven supervision, while the underlying VLM remains frozen.
TABLE II: Datasets used in our experiments and their coverage of hallucination types.
Dataset #Samples Split Ground-Truth Verification Object Relation / Scene False Premise
HalLoc-VQA 80,540 / 10,932 / 10,082 Train/Val/Test Token-level verification
POPE 9,000 Test Object-level verification
HaloQuest 7,140 / 608 Train/Test Answer-level verification Partial
VQA v2 443,757 / 214,354 / 447,793 Train/Val/Test Answer-level verification
HalLoc overall contains about 155k labeled samples across tasks including VQA; the VQA subset used in our work is drawn from this set.
V Experimental Setup
V-A Datasets
We evaluate our approach on four multimodal datasets with complementary characteristics: HalLoc-VQA [park2025halloc], POPE [li-etal-2023-evaluating], HaloQuest [wang2024haloquest], and VQA v2 [goyal2017vqa], which differ in hallucination types, annotation granularity, and inducing mechanisms. HalLoc-VQA provides token-level supervision for object- and scene-level hallucinations. POPE focuses on object-level hallucinations via direct probing of object existence, HaloQuest emphasizes false-premise and reasoning-induced hallucinations, and VQA v2 serves as a naturalistic evaluation set for real-world visual grounding challenges. Dataset statistics are summarized in Table II; for training and evaluation, we sample 1,000 instances from each split (full HaloQuest test set used) and report averages over three random seeds, ensuring consistent evaluation while controlling computational cost, following common practice in multimodal hallucination and uncertainty analysis [lee2024vhelm, chen-etal-2025-unveiling, park2025glsim, farquhar2024detecting].
V-B Baselines
We compare FaithSCAN with a diverse set of representative baselines for multimodal hallucination detection, focusing on uncertainty-based methods at token, latent, and semantic levels. Token-level baselines include Predictive Entropy (PE) [holtzman2020curious] and Token-level Negative Log-Likelihood (T-NLL) [zhang2023enhancing], measuring per-token predictive uncertainty. Latent-space baselines such as Embedding Variance (EmbVar) [grewal2024improving] and Semantic Embedding Uncertainty (SEU) [grewal2024improving] capture instability in vision-conditioned answer embeddings. Semantic-level baselines include Semantic Entropy (SE) [farquhar2024detecting] and a multimodal adaptation of SelfCheckGPT [manakul2023selfcheckgpt], detecting contradictions across answer pairs. A multimodal $P(\text{True})$ verification baseline [farquhar2024detecting] evaluates answer correctness via image-conditioned yes/no prompts. Finally, a Logistic Regression [hosmer2013applied] classifier trained on our dataset serves as a supervised reference.
V-C Implementation Details and Metrics
Experiments are conducted on three vision–language models: IB-T5-XL, LLaVA-8B, and Qwen3-VL-8B [^1] Full model names: IB-T5-XL (Salesforce/instructblip-flan-t5-xl); LLaVA-8B (llava-hf/llama3-llava-next-8b-hf); Qwen3-VL-8B (Qwen/Qwen3-VL-8B-Instruct). . Uncertainty signals, including token-level log-likelihood, predictive entropy, and hidden representations from a selected transformer layer (see Section VI-C3 for layer-wise analysis), are extracted in a single forward pass during decoding. Hallucination supervision is derived via a model-driven verification scheme using Qwen2.5-VL-32B-Instruct, with labels obtained through low-temperature sampling (temperature 0.1) to validate generated answers. Each uncertainty branch is projected to a 64-dimensional latent space and fused via a lightweight attention module. Models are trained for 20 epochs using AdamW, with learning rate $1\times 10^{-4}$ and batch size 32, employing early stopping on 2×RTX 3090 GPUs, and inference is performed on a single GPU.
Focusing on the I2T setting where evaluation is performed primarily on generated textual answers, we assess hallucination detection using AUROC (ROC) to measure overall ranking quality, AURAC (RAC) to evaluate selective prediction under rejection, F1@Best (F1@B) to report the optimal classification trade-off, and RejAcc@50% (RA@50) to quantify accuracy after rejecting the most uncertain half of predictions [farquhar2024detecting, zhao2025sese, tong2025semantic]. Together, these metrics provide a comprehensive and complementary evaluation of hallucination detection performance from both ranking and decision-making perspectives.
TABLE III: RQ1: Overall ID performance on three datasets and three backbone models. Latency is reported as the average sampling and inference time per 100 examples (in seconds), with colors indicating increasing computational cost from green to red. In the Sample column, red marks the slowest methods requiring multiple samplings per query. Bold and underlined values denote the best and second-best results, respectively.
HalLoc-VQA HaloQuest VQA v2 Latency /100 (s)
ROC RAC F1@B RA@50 ROC RAC F1@B RA@50 ROC RAC F1@B RA@50 Sample Infer
PE 0.536 0.526 0.730 0.527 0.578 0.526 0.805 0.534 0.624 0.239 0.561 0.225 1326.308 0.002
T-NLL 0.535 0.537 0.762 0.542 0.571 0.550 0.730 0.549 0.623 0.265 0.502 0.285 1257.512 0.001
EmbVar 0.657 0.464 0.746 0.453 0.605 0.521 0.790 0.578 0.724 0.773 0.667 0.845 1430.884 0.060
SE 0.583 0.563 0.653 0.562 0.537 0.595 0.805 0.608 0.565 0.311 0.491 0.297 1381.929 4.913
SEU 0.663 0.452 0.738 0.440 0.660 0.498 0.776 0.515 0.772 0.787 0.667 0.868 1685.975 1.845
SelfCheckGPT 0.509 0.545 0.722 0.590 0.474 0.623 0.797 0.637 0.502 0.684 0.514 0.657 1602.887 78.613
$P(\text{True})$ 0.675 0.597 0.753 0.565 0.515 0.609 0.770 0.642 0.703 0.782 0.554 0.777 1127.961 14.938
Logistic 0.662 0.657 0.728 0.666 0.646 0.313 0.858 0.332 0.733 0.740 0.580 0.810 1570.349 0.087
IB-T5-XL FaithSCAN (Ours) 0.778 0.824 0.778 0.860 0.709 0.665 0.684 0.675 0.786 0.813 0.671 0.810 1560.237 0.605
PE 0.605 0.276 0.535 0.258 0.440 0.368 0.515 0.348 0.553 0.588 0.433 0.565 1491.448 0.002
T-NLL 0.607 0.256 0.529 0.247 0.447 0.372 0.446 0.373 0.558 0.725 0.413 0.787 1524.740 0.001
EmbVar 0.747 0.820 0.509 0.855 0.546 0.355 0.515 0.284 0.593 0.437 0.481 0.345 2504.584 0.175
SE 0.683 0.747 0.462 0.795 0.595 0.280 0.557 0.275 0.630 0.186 0.425 0.172 2632.833 11.212
SEU 0.769 0.799 0.667 0.865 0.536 0.324 0.434 0.328 0.604 0.666 0.429 0.645 2443.890 1.769
SelfCheckGPT 0.715 0.779 0.625 0.818 0.603 0.576 0.510 0.534 0.693 0.741 0.533 0.755 2588.092 340.683
$P(\text{True})$ 0.615 0.690 0.520 0.748 0.716 0.761 0.553 0.730 0.663 0.772 0.564 0.825 1178.084 19.511
Logistic 0.719 0.771 0.566 0.800 0.792 0.838 0.729 0.857 0.633 0.786 0.432 0.796 1874.895 0.114
LLaVA-8B FaithSCAN (Ours) 0.759 0.801 0.643 0.856 0.681 0.773 0.600 0.762 0.733 0.848 0.536 0.850 1898.950 0.741
PE 0.686 0.178 0.535 0.200 0.573 0.726 0.229 0.789 0.549 0.394 0.428 0.305 1328.434 0.002
T-NLL 0.650 0.233 0.524 0.190 0.542 0.733 0.257 0.784 0.524 0.193 0.312 0.195 1330.996 0.001
EmbVar 0.628 0.705 0.570 0.787 0.641 0.875 0.444 0.902 0.587 0.777 0.483 0.823 2194.586 0.107
SE 0.559 0.325 0.456 0.305 0.606 0.873 0.316 0.873 0.532 0.208 0.374 0.205 2019.877 13.812
SEU 0.653 0.742 0.587 0.800 0.678 0.868 0.500 0.917 0.614 0.805 0.396 0.848 2199.427 1.821
SelfCheckGPT 0.656 0.716 0.552 0.748 0.783 0.878 0.501 0.903 0.712 0.851 0.430 0.877 2165.013 367.522
$P(\text{True})$ 0.750 0.825 0.548 0.838 0.631 0.776 0.394 0.832 0.777 0.860 0.499 0.915 1143.720 24.299
Logistic 0.656 0.448 0.493 0.440 0.746 0.294 0.422 0.254 0.610 0.438 0.354 0.416 1689.458 0.102
Qwen3-VL-8B FaithSCAN (Ours) 0.810 0.864 0.671 0.860 0.742 0.886 0.505 0.921 0.738 0.885 0.510 0.920 1741.652 1.191
VI Experiments
In this section, we conduct extensive experiments to evaluate the proposed method from multiple perspectives. Specifically, our experimental study is organized around four research questions. RQ1 (Overall Effectiveness and Generalization): How effective is the proposed method in detecting hallucinations across different vision–language models and datasets covering diverse hallucination types? RQ2 (Model-driven Supervision Analysis): How effective and reliable is the proposed model-driven verification strategy in providing supervision signals for hallucination detection? RQ3 (Internal Representation Analysis): What roles do different internal representations, including token-level uncertainty, visual embeddings, and aligned visual–semantic embeddings, play in hallucination detection, and how do they complement each other? RQ4 (Practical Utility and Interpretability): How does the proposed method perform under practical decision-making scenarios, including its behavior at different recall levels and the interpretability of its internal uncertainty signals?
VI-A Overall Effectiveness and Generalization (RQ1)
We evaluate FaithSCAN under in-distribution (ID) and out-of-distribution (OOD) settings across four datasets, using 1,000 randomly sampled instances per split and averaging results over three runs. FaithSCAN leverages token-level uncertainty signals (log-likelihood, entropy, embeddings) and visual patch features, with sequence-level representations obtained via global average pooling and fused through attention (see Section VI-C for architectural choices).
VI-A1 In-distribution Evaluation
ID results (Table III) show that FaithSCAN consistently outperforms strong baselines, achieving up to 10% AUROC improvement and maintaining top performance across metrics and backbone models. Sampling-based baselines such as SelfCheckGPT and $P(\text{True})$ perform competitively but incur higher inference costs, whereas FaithSCAN achieves strong detection in a single pass. SE underperforms due to long-form, detailed answers that undermine clustering assumptions. Performance further improves with more capable backbones, suggesting richer internal representations provide clearer uncertainty signals.
TABLE IV: RQ1: OOD hallucination detection performance (AUROC) under cross-dataset and multi-source training. In cross-dataset evaluation, models are trained on a single dataset and tested on unseen datasets, while multi-source results are obtained by joint training on all available datasets. L and F denote logistic regression and FaithSCAN (ours), respectively. AUROC values below 0.5 are highlighted in red, indicating generalization failure, and for multi-source training, the better result between L and H is highlighted in green.
Train $\rightarrow$ Test IB-T5-XL LLaVA-8B Qwen3-VL-8B
L F L F L F
Single-source training (cross-dataset evaluation)
HalLoc-VQA $\rightarrow$ POPE 0.499 0.603 0.671 0.703 0.559 0.632
HalLoc-VQA $\rightarrow$ HaloQuest 0.514 0.508 0.601 0.393 0.493 0.412
HalLoc-VQA $\rightarrow$ VQA v2 0.631 0.669 0.546 0.571 0.578 0.624
HaloQuest $\rightarrow$ HalLoc-VQA 0.555 0.567 0.501 0.506 0.521 0.429
HaloQuest $\rightarrow$ POPE 0.641 0.573 0.580 0.511 0.631 0.639
HaloQuest $\rightarrow$ VQA v2 0.569 0.561 0.567 0.585 0.579 0.594
VQA v2 $\rightarrow$ HalLoc-VQA 0.628 0.640 0.598 0.571 0.556 0.605
VQA v2 $\rightarrow$ POPE 0.629 0.712 0.628 0.604 0.644 0.701
VQA v2 $\rightarrow$ HaloQuest 0.541 0.664 0.700 0.759 0.650 0.717
Multi-source training (trained on all datasets)
All $\rightarrow$ HalLoc-VQA 0.703 0.743 0.663 0.700 0.611 0.672
All $\rightarrow$ HaloQuest 0.696 0.697 0.777 0.796 0.757 0.749
All $\rightarrow$ VQA v2 0.709 0.715 0.631 0.638 0.618 0.668
All $\rightarrow$ POPE 0.680 0.744 0.669 0.675 0.654 0.689
Moreover, we evaluated the computational efficiency of each method in both the sampling and inference stages. In the sampling phase (e.g., obtaining answers or extracting embeddings), most existing methods are relatively slow. Regarding inference, approaches that rely on multiple samples or use external models exhibit far lower efficiency than those that infer directly from the model’s internal states. Because FaithSCAN integrates multiple embeddings, its inference speed is slightly reduced when handling larger embedding sizes; nevertheless, it remains among the fastest, highlighting the advantages of our approach.
VI-A2 Out-of-distribution Evaluation
OOD evaluation (Table IV) reveals limited generalization for models trained on a single dataset, especially when hallucination types differ (e.g., object/scene-level vs. false-premise). FaithSCAN exhibits stronger transfer than logistic regression, and multi-source training enhances OOD performance, highlighting the importance of diverse training data. Nonetheless, a performance gap between ID and OOD underscores the challenge of hallucination detection under distribution shifts, motivating future work on improving generalization of model-driven uncertainty approaches.
Figure 3: RQ2: Illustrative examples of the consistency judgment process based on model-driven supervision, showing both positive and negative cases. (a) The sample is drawn from POPE, where the LLM-based judger correctly identifies a contradiction relationship between the model-generated answer and the ground-truth reference. (b) The sample is drawn from VQA v2, where the generated answer is not strictly precise; nevertheless, it is judged as entailment by the LLM judger.
VI-B Model-driven Supervision Analysis (RQ2)
We evaluate whether the proposed model-driven supervision reliably approximates human annotations for hallucination detection.
Qualitative examples are shown in Fig. 3. In Fig. 3a, a POPE sample demonstrates correct detection of a contradiction by the LLM-based judger, reflecting the dataset’s clear object-level hallucination boundaries. Fig. 3b illustrates a failure case from VQA v2, where minor factual errors are masked by semantic similarity, leading to an incorrect entailment judgment.
TABLE V: RQ2: Consistency analysis between model-driven supervision signals and ground-truth hallucination annotations across different datasets.
Dataset Agreement $\uparrow$ Cohen’s Kappa $\uparrow$ MCC $\uparrow$
HalLoc-VQA 0.960 0.834 0.846
POPE 0.990 0.872 0.879
HaloQuest 0.960 0.816 0.829
VQA v2 0.980 0.845 0.852
Quantitatively (Table V), the supervision achieves high agreement with human labels (0.960–0.990), with Cohen’s Kappa and MCC exceeding 0.81 across datasets. POPE shows the highest consistency due to its simpler object-level hallucinations, while more complex datasets like HalLoc-VQA, HaloQuest, and VQA v2 yield slightly lower but still strong agreement.
These results indicate that the model-driven supervision effectively aligns with human judgment, providing a reliable signal for hallucination detection across diverse scenarios.
Figure 4: RQ3: AUROC scores for different uncertainty source combinations used in hallucination detection across four datasets and three base models. Labels indicate uncertainty types: T = text-level (log-likelihood, entropy, embedding), P = patch-level visual, and A = cross-modal alignment.
(a) Results of IB-T5-XL.
(b) Results of LLaVA-8B.
(c) Results of Qwen3-VL-8B.
Figure 5: RQ3: Ablation study on feature combination, sequence modeling strategy, and fusion methods on the HalLoc-VQA training set. The fusion methods include: Global Avg (average of all token-level features), CNN Emb (CNN modeling only on the embeddings), and CNN All (CNN modeling applied to all features).
VI-C Internal Representation Analysis (RQ3)
We investigate why FaithSCAN effectively detects hallucinations by analyzing the role of different internal representations and uncertainty sources. Experiments are conducted on HalLoc-VQA, POPE, HaloQuest, and VQA v2 using IB-T5-XL, LLaVA-8B, and Qwen3-8B.
VI-C1 Contribution of Independent Uncertainty Sources
We investigate the contributions of independent uncertainty sources derived from distinct internal representations, including token-level uncertainty from language generation, visual uncertainty from compressed patch embeddings, and visual–semantic uncertainty from features aligned to the textual semantic space.
We first evaluate each source in isolation. As shown in Fig. 4, visual patch–level and visual–semantic uncertainty alone generally yield limited discriminative power across datasets, indicating that hallucinations are difficult to detect without access to token-level generation signals. In contrast, token-level uncertainty consistently provides a strong baseline, suggesting that many hallucination phenomena are rooted in the internal states of the language generation process.
The role of visual uncertainty, however, varies substantially across datasets and hallucination types. On POPE, which predominantly focuses on object-level hallucinations, integrating pure visual patch uncertainty leads to significant performance gains. This result suggests that object hallucinations are closely tied to perceptual failures, and that uncertainty at the visual perception level constitutes a critical failure pathway in such cases.
In contrast, on HalLoc-VQA and VQA v2, incorporating visual features yields limited or inconsistent improvements. These datasets involve more complex reasoning and compositional question answering, where hallucinations are less attributable to visual ambiguity and more strongly associated with failures in language modeling and multimodal reasoning. This observation reflects a shift in the dominant failure pathway from perception to higher-level reasoning processes.
A similar pattern is observed on HaloQuest, which primarily targets false-premise hallucinations. For models with relatively weaker visual capabilities, incorporating visual uncertainty substantially improves detection performance, indicating that interaction-level hallucinations can originate from unreliable visual grounding. However, for models with stronger visual encoders, the gains diminish, suggesting that hallucinations in this setting are more likely driven by reasoning-level failures. This further indicates that the source of hallucination is architecture-dependent and cannot be attributed to a single modality.
Overall, these results empirically support our methodological premise: effective hallucination detection requires explicitly modeling the mapping between observed hallucination phenomena, their manifestations in VQA, the underlying multimodal failure pathways, and the corresponding internal model states.
VI-C2 Ablation Study on Token-level Uncertainty Modeling
We evaluate token-level uncertainty signals derived from log-likelihoods, entropies, and hidden embeddings. As shown in Fig. 5, embeddings provide the richest signal, followed by entropy, while log-likelihood contributes marginally. Combining these signals consistently improves performance across datasets and models, with attention-based fusion outperforming simple concatenation. Incorporating CNN-based sequence modeling on embeddings further enhances detection, highlighting that localized anomalies in embeddings capture hallucinations effectively. Based on the results in Fig. 5, we adopt the combination of log-likelihood, entropy, embeddings, global average pooling, and attention as a stable FaithSCAN variant, which is used in all subsequent experiments.
Figure 6: RQ3: Layer-wise contribution analysis of hallucination detection performance. The best-performing layer is highlighted, and $\Delta$ vs Mean denotes the relative improvement of the optimal layer over the average performance.
VI-C3 Layer-wise Representation Analysis
We evaluated token-level representations from different decoder layers for hallucination detection on HalLoc-VQA using IB-T5-XL, LLaVA-8B, and Qwen3-8B. Each layer was tested on 100 samples, repeated three times, using all uncertainty signals from Section IV-D with attention-based fusion. Fig. 6 shows that representations from the earliest and deepest layers consistently underperform, while intermediate layers achieve higher performance across all models. The layer-wise uncertainty exhibits fluctuations, indicating that different layers encode complementary information: intermediate layers balance token-level variability and semantic abstraction, capturing cross-modal cues relevant to hallucination. This trend is particularly clear for Qwen3-8B and LLaVA-8B.
Based on these results, we select intermediate layers as token-level features in subsequent experiments. For most models we use the 2/3 depth layer, while for InstructBLIP we choose the 1/2 depth layer. This choice is guided by the layer-wise performance observed in this analysis and balances predictive performance with computational efficiency.
Figure 7: RQ4: ROC and PR curves of FaithSCAN and baseline methods on the HalLoc-VQA $\times$ Qwen3-VL-8B setting.
VI-D Practical Utility and Interpretability (RQ4)
Hallucination detectors are typically deployed by applying a threshold on uncertainty or confidence scores, making performance sensitive to the precision–recall trade-off. We evaluate FaithSCAN under the HalLoc-VQA $\times$ Qwen3-VL-8B setting, reporting ROC and PR curves alongside baselines, and visualize normalized uncertainty distributions to illustrate separation between hallucinated and non-hallucinated samples.
To assess interpretability, we analyze token-level features using a gradient-based approach inspired by saliency analysis [simonyan2014deep, shrikumar2017learning]. For input features $\mathbf{x}={x_{t,f}}$ and predicted probability $\hat{y}=\sigma(f_{\theta}(\mathbf{x}))$, attributions are computed as $A_{t,f}=(\partial\hat{y}/\partial x_{t,f})\cdot x_{t,f}$, capturing both a feature’s influence on the prediction and its signal strength, yielding interpretable token-level contributions.
Figure 8: RQ4: Violin plots of normalized uncertainty scores for hallucinated and non-hallucinated samples.
Figure 9: RQ4: Token-level feature attribution visualization using the grad $\times$ input method. Positive contributions highlight hallucination-related tokens, while negative contributions correspond to non-hallucinated regions.
Fig. 7 presents the ROC and PR curves of FaithSCAN and competing baselines. Across most operating points, FaithSCAN consistently achieves superior performance, maintaining higher precision under the same recall levels.
To further examine the uncertainty modeling behavior, Fig. 8 visualizes the normalized uncertainty score distributions using violin plots. Compared to baseline methods, FaithSCAN produces a markedly larger separation between hallucinated and non-hallucinated samples. Notably, although FaithSCAN is trained using binary supervision, it implicitly learns a meaningful uncertainty score that correlates well with hallucination likelihood.
Finally, Fig. 9 provides a qualitative case study using token-level attributions obtained via the grad $\times$ input method. Due to the difficulty of aligning visual features with textual tokens, we focus on text-side feature channels in this analysis. In the illustrated example, Qwen3-VL-8B incorrectly identifies a stool instead of a chair. The key semantic distinction lies in the presence of a backrest, which is overlooked by the model despite rich visual details. FaithSCAN assigns strong positive contributions to the corresponding hallucinated token sequence, accurately capturing the erroneous reasoning process. These highlighted regions align well with human intuition, demonstrating the interpretability and diagnostic value of our approach.
VII Conclusion
This paper investigates faithfulness hallucination detection in VQA from a model internal perspective. Our results indicate that many hallucinations are already encoded in the internal states of VLMs during generation, rather than emerging solely at the output level. Uncertainty signals from language decoding, visual perception, and cross modal interaction provide complementary views of different hallucination failure pathways, and their relative contributions vary across model architectures and hallucination types.
More broadly, this work suggests that hallucinations in VQA are structured phenomena rooted in multimodal reasoning dynamics. The effectiveness of lightweight and trainable detectors operating on frozen model states demonstrates that reliable hallucination detection can be achieved without repeated sampling or external verification. We hope these findings motivate future research to treat hallucination as an intrinsic property of vision language reasoning, supporting the development of more reliable and interpretable multimodal systems.